OpenStudy (anonymous):

How do you convert two points to standard form(ax+by=c)? points are 1,2 and 4,4 please include explalation

OpenStudy (ranga):

You want the equation of the line passing through the points (1,2) and (4,4) in standard form. Slope of the line m = (y2 - y1) / (x2 - x1) = (4-2)/(4-1) = 2/3 Equation of the line is: (y-y1) = m(x-x1) y - 2 = 2/3(x - 1) multiply both sides by 3 3y - 6 = 2x - 2 2x - 3y = -6 + 2 2x - 3y = -4 is the equation of the line in standard form.

OpenStudy (anonymous):

@ranga wont we distruble a -1 one to make the -4 postive

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Some follow the convention to make the coefficient of x positive others make c positive. Depends on what your textbook/teacher say.

OpenStudy (anonymous):

ok thanks